### v1

# Leveraging non-unital noise for gate-based quantum reservoir computing

Abstract
Despite theoretical promises, most machine learning algorithms fail on current quantum computers due to overwhelming noise. Nevertheless, many recurrent architectures, are most efficient if they can process information by progressively reducing the emphasis on inputs from earlier time steps. We identify a noise model that ensures the functioning of an echo state network employing a gate-based quantum computer for reservoir computing applications. Energy dissipation induced by amplitude damping drastically improves the short-term memory capacity and expressivity of the network, by simultaneously providing fading memory and richer dynamics. There is an ideal dissipation rate that ensures the best operation of the echo state network around $\gamma\sim$ 0.03. Nevertheless, these beneficial effects are stable as the intensity of the applied noise increases. The improvement of the learning is confirmed by emulating a realistic noise model applied to superconducting qubits, paving the way for the application of reservoir computing methods in current non-fault-tolerant quantum computers.
I Introduction
The intrinsic noise in noisy intermediate-scale quantum (NISQ) computers poses a major limitation in many machine-learning protocols [1], restricting the natural application of quantum computing in artificial intelligence [2, 3]. Nevertheless, dissipation may serve as a fully-fledged resource for quantum computation [4]. Our results show that the action of specific noise, namely noise modeled by non-unital channels, drastically improves the short-term memory capacity and expressivity of a quantum network, proving the reliability of quantum reservoir computing as a valid computational architecture for memory-dependent tasks employing a gate-based quantum computer. We emphasize that an in-depth study of the beneficial computational effects of non-unital noise in quantum machine learning is only in its early stages [5], unveiling, for example, a promising application in avoiding barren plateaus in variational problems [6].
Reservoir computing is a well-established supervised machine learning algorithm employing a fixed neural network, the reservoir, to process time-dependent information. In addition to initial digital and neuromorphic proposals, respectively, the echo state network [7] and the liquid state machine [8], reservoir computing has proven effective in a wide range of unconventional deployments [9]. Among several physical implementations [10], the exponentially large computational capacity of quantum systems has been harnessed for reservoir computing [11, 12], employing photonic circuits [13, 14, 15], bosonic oscillators [16, 17], fermionic systems [18, 19], neutral atoms [20] and spin networks [21, 22, 23, 24, 25, 26]. Focusing on superconductive quantum computers, we refer to Ref. [27, 28] for early implementations, more recently empowered by mid-circuit measurements [29].
Reservoir computing appears particularly suitable for leveraging quantum computing, as it does not require any training of the parameters of the quantum evolution. An essential feature essential for reservoir computing is the so-called fading memory [30, 31]. Namely, a reservoir network is most efficient if the information is processed by progressively reducing the emphasis on inputs from earlier time steps. Indeed, dissipation has recently been recognized as a powerful computational resource in reservoir architectures employing spin dynamics [25]. Regarding superconductive quantum computers, artificial manual reset of the information flow has proven to be a reliable resource for fading memory in memory-dependent tasks [28]. Here, we step further, by demonstrating how the intrinsic noise in superconducting quantum computers can be harnessed to spontaneously ensure the functioning of a quantum echo state network.
By testing our quantum echo state network on standard memory-dependent benchmarks, we recognize which type of noise has beneficial effects on the performances, distinguishing it from detrimental noise. Thus, we show that noise modeled by a non-unital channel is needed to simultaneously guarantee memory capacity and accuracy in nonlinear time-dependent tasks, confirming early results for memory-independent classification tasks [32]. Remarkably, the quantum channel emulating realistic noise on a superconductive quantum computer falls within this class. As typically observed in recurrent networks [33, 34], we identify a critical regime tuned by noise intensity, at which several network capabilities, such as short-term memory capacity and expressivity, are maximized. Interestingly such computational capabilities are preserved under a considerable increase in noise intensity if a complete measurement of the circuit is performed, suggesting immediate possible real-world application on NISQ superconductive computers. Finally, we investigate the effect of local measurements, since it is known that the measurement rate drastically influences the emerging properties of a register of qubits [35]. While the suboptimal regime is rapidly disrupted by increasing noise intensity, the optimal one persists with almost unchanged performances, allowing for a reduction of required measurements by employing error control protocols. Our proof-of-concept paves the way for real-world application of quantum reservoir computing on both noisy intermediate-scale and early fault-tolerant quantum computers.
Figure 1: The architecture of the quantum echo state network. Each layer of unitary gates is the composition of rotations along $X$ and $Z$ axis of the Bloch sphere and CNOT operators so that the qubits of the register are entangled.
II Methods
II.1 Quantum reservoir computing and fading memory
Denoting $\mathcal{S}(\mathcal{H})$ the set of density operators on the Hilbert space $\mathcal{H}$ that describes some quantum system, a quantum reservoir computer is described by the equations
$\begin{cases}\rho_{t+1}=\mathcal{T}(\rho_{t},u_{t+1})\\ y_{t+1}=h(\rho_{t+1})\end{cases}$
where $\mathcal{I}$, is the space of classical input sequences, $\mathcal{I}\ni u=\{\dots,u_{-1},u_{0}\}$ is the input signal and $\mathcal{T}:\mathcal{S}(\mathcal{H})\rightarrow\mathcal{S}(\mathcal{H})$ is a quantum channel. The key property of a reservoir computer is the well-known echo state property (or, convergence property) [36], which ensures that for each input sequence $u$, there exists a unique reservoir signal $z$. If the echo state property holds, the reservoir computer defines a unique, well-posed operator (or, filter in the language of reservoir computing) defined by
$T(u)_{t}=\mathcal{T}(T(\rho)_{t-1},u_{t})$
with $u\in\mathcal{I}$, and then a unique functional on the set of input sequences $T:\mathcal{I}\rightarrow\mathcal{S}(\mathcal{H})$, defined as $T(u)=T(u)_{0}$. Note that the choice of $t=0$ as the final time step is merely irrelevant since the system is time-invariant. Fading memory is formally defined as the continuity of this functional, and thus of related filter since the relation is bijective, with respect to a proper topology.
Definition II.1 (Fading memory)
Let $\omega\colon\mathbb{Z}^{-}\rightarrow(0,1]$ be an increasing function with zero limit at infinity and $\omega(0)=1$ and define the distance
$d_{\omega}(u,v)=\sup_{t\in\mathbb{Z}^{-}}\norm{u_{t}-v_{t}}\omega_{t},\quad\forall u,v\in\mathcal{I}\,.$
A reservoir computer has fading memory if the associated functional $T:\mathcal{(}I)\rightarrow\mathbb{R}$ is continuous in the topology of the metric space $\left(\mathcal{I},d_{\omega}\right)$.
Fading memory of a quantum reservoir ultimately depends on the properties of the quantum channel $\mathcal{T}$ that defines the quantum reservoir computer. In particular, we have the following important result, that ensures fading memory.
Theorem II.2 (Theorem 3 in [37])
If the quantum channel $T$ is strictly contractive in the trace norm $\norm{A}_{1}=\mathrm{tr}\left(\sqrt{A^{\dagger}A}\right)$, namely
$\norm{\mathcal{T}(\rho)-\mathcal{T}(\sigma)}_{1}\leq r\norm{\rho-\sigma}_{1}$
for some $r\in(0,1)$, then the associated reservoir computer has the echo state property and fading memory.
Remark II.3
Any quantum channel in non-expansive in trace norm, namely $\norm{\mathcal{T}(\rho)}_{1}\leq\norm{\rho}_{1}$.
II.2 Architecture of the quantum echo state network
Our quantum echo state network exploits a $N$-qubits quantum register to encode and process time-dependent information and reproduce an input-output nonlinear mapping $S(u)=\hat{y}$ between two time-dependent sequences $u=\left\{u_{t}\right\}_{t=0,\dots,L}$ and $\hat{y}=\left\{\hat{y}_{t}\right\}_{t=0,\dots,L}$. We refer to $u$ and $\hat{y}$ as the input and the target sequence, respectively.
Denoting with $\rho_{t}$ the density operator describing the $N$-qubits system at time $t$, its temporal evolution $\rho_{t+1}=\mathcal{T}_{u_{t+1}}(\rho_{t})$ is the composition of a noisy quantum channel, including the unitary gates that encode the input signal $u$, and a mid-circuit measurement operation. Precisely, at each time step, the system first evolves with the quantum evolution
$\widetilde{\rho}_{t+1}=\mathcal{E}\left(U(u_{t+1})\rho_{t}U^{{\dagger}}(u_{t+1})\right)\,.$ (1)
In Equation 1, $U(u_{t+1})$ represents some unitary gates implemented by the quantum computer. Our model includes CX operators and rotations, whose angles are dependent on the incoming input value. The precise architecture adopted in our work is described in Figure 1. $\mathcal{E}$ is a quantum channel that models the noise in the quantum register. We will specify the noise models considered in this work later on. Then, a mid-circuit measurement of the register is performed and the system is let evolve in the collapsed state. Formally, we have
$\rho_{t+1}=\Pi_{M}\widetilde{\rho}_{t+1}\Pi_{M}^{\dagger}$
with, denoting with $m_{s,t+1}\in\left\{0,1\right\}$ the outcome of the measure of the qubits $s$ at time $t+1$,
$\Pi_{M}=\bigotimes_{s=1}^{M}|m_{s,t+1}\rangle\langle m_{s,t+1}|\otimes\mathbb{I}_{N-M}$
Here, $M$ is the number of measured qubits in the register at each time step. We remark that both the case of a complete measurement, $M=N$, and the case of a partial measurement, $M<N$, are treated in the Results. The measurement results are taken as the reservoir signal passes throughout the processing. Explicitly, the reservoir signal at time $t$ is the expectation value of the $Z$ Pauli operator of each qubit, taken separately,
$z_{t}=\left[\langle Z_{1}\rangle_{\rho_{t}},\dots,\langle Z_{M}\rangle_{\rho_{t}}\right]^{T}\,.$
In our numerical experiments, the expectation value of each $Z_{i}$ is obtained by repeating the experiment for $S\sim 10^{5}$ shots and computing $\langle Z_{i}\rangle_{\rho_{t}}=\frac{1}{S}\left(\sum 1_{m_{i,t}=0}-\sum 1_{m_{i,t}=1}\right),$ where the sum is taken over all shots.
II.3 Training of linear readout
Reservoir computing aims to reproduce best a given nonlinear map $S(u)=\hat{y}$ by training the linear readout $h$. We will give examples of such learning tasks $S(u)$ later in the work. After an initial washout period $T_{\text{wo}}$, required to erase the dependence on the initial state of the network, the reservoir signal is extracted until the end of the training interval $T_{\text{tr}}$ and it is stored in the matrix $H=\left\{z_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$. After the quantum evolution, the optimal readout weights are computed. Precisely, the weights of the linear readout $y=Hw$ are chosen as the set of parameters $w$ that minimizes the distance $\norm{\hat{y}-y}$. It can be easily computed by linear regression or exploiting the Moore-Penrose pseudoinverse matrix, namely $w=(H^{T}H)^{-1}H^{T}\hat{y}_{\text{tr}}$, where $\hat{y}_{\text{tr}}=\left\{\hat{y}_{t}\right\}_{t=T_{\text{wo}}+1,\dots,T_{\text{tr}}}$ is the portion of target sequence used for the training. We refer to Reference [28] for further details on the training algorithm. After the training, the predicted values of $y$ are given by
$y_{t}=z_{t}\cdot w\qquad\text{for any }t\in[T_{tr}+1,L]\,.$
It is worthwhile to notice that, in the exact spirit of reservoir computing, the same reservoir signal can be exploited to reproduce different maps, only repeating the training to compute the optimal weights. This fact allows for the reuse of information stored in the reservoir for different tasks that share the same input, drastically reducing execution time.
Table 1: The Kraus decomposition of the quantum channel describing respectively amplitude damping, phase damping and depolarizing noise.
Noise model Description Kraus decomposition
Amplitude damping Loss of energy to the environment $K_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix},\,\,K_{1}=\begin{pmatrix}0&\sqrt{\gamma}\\ 0&0\end{pmatrix}$
Phase damping Loss of quantum information $K_{0}=\sqrt{1-\gamma}\,\mathbb{I},\,\,K_{1}=\begin{pmatrix}\sqrt{\gamma}&0\\ 0&0\end{pmatrix},$ $K_{2}=\begin{pmatrix}0&0\\ 0&\sqrt{\gamma}\end{pmatrix}$
Depolarizing Decay to maximally mixed state $K_{0}=\sqrt{1-\gamma}\,\mathbb{I},\,\,K_{1}=\sqrt{\gamma}\,X$ $K_{2}=\sqrt{\gamma}\,Y,\,\,K_{3}=\sqrt{\gamma}\,Z$
II.4 Performance analysis
Our experiments aim to demonstrate the beneficial effect of intrinsic noise in a superconductive quantum computer for reservoir computing. First, to verify whether quantum noise improves the fading memory, we study its effect on the correlation among different measurement outcomes of the same qubit over time.
Then, the performances of our computational architecture under the effect of quantum noise are analyzed by exploiting standard benchmarks for recurrent neural networks. We measure the short-term memory capacity of our network using the memory capacity metric, which quantifies the amount of variance in the delayed input that can be recovered from the trained output [7]. To compute it, the system is required to replicate an input sequence $\left\{u_{t}\right\}_{t=0}^{T}$ delayed by a time interval $d$. Namely, the system is required to approximate the functional $\hat{y}_{t}=u_{t-d}$. Then, the $d$-memory capacity is defined as
$\mathrm{MC}_{d}=\frac{\mathrm{Cov}\left(y,\hat{y}\right)}{\mathrm{Var}(y)\mathrm{Var}(\hat{y})}\,,$
where $y$ is the vector of the predicted values. The memory capacity of the network is calculated as the sum up to a certain $d_{\text{max}}$,
$\mathrm{MC}=\frac{1}{d_{\text{max}}}\sum_{d=1}^{d_{\text{max}}}\mathrm{MC}_{d}\,,$
normalized so that $\mathrm{MC}\in[0,1]$. We remark that $\mathrm{MC}$ values closer to 1 indicate a higher memory capacity of the system. We fix $d_{\text{max}}=10$ in the numerical experiments.
The predictive capabilities of our echo state network are tested with the so-called NARMA task. It is a nonlinear filter with past dependence, commonly used as a benchmark for the computational power of time-dependent learning systems [38]. Precisely, the NARMA-$p$ task is formally defined as
$\hat{y}_{t}=0.4y_{t}+0.1\left(\sum_{l=0}^{p-1}y_{t-l}\right)+u_{t-p+1}+0.1\,.$
We pick $p=2,5\text{ and }8$ in the paper. The accuracy of the prediction is quantified with the normalized mean-square error, expressed as
$\mathrm{NSME}(y,\hat{y})=\frac{\sum_{t=T_{\text{tr}+1}}^{t=L}|\hat{y}_{t}-y_{t}|^{2}}{\sum_{t=T_{\text{tr}+1}}^{t=L}\hat{y}_{t}^{2}}$
where $y$ is the vector of the predicted values after training.
II.5 Ideal and realistic quantum noise models
The effect of noise in a quantum computer can be simulated by repeatedly applying a quantum channel to the density operator that describes the state of the qubit register. Precisely, in numerical simulations, a quantum channel $\mathcal{E}$, representing a certain noise model, is applied to each qubit after the action of each unitary gate $U$, as shown in Equation  1. First, we investigate the effect of three typical ideal models of quantum noise. Namely, we test the effect of amplitude damping, phase damping, and depolarizing noises. We recall that a quantum channel can always be expressed in terms of its Kraus decomposition,
$\mathcal{E}(\rho)=\sum_{i=1}^{r}K_{i}\rho K_{i}^{\dagger}\,.$
We refer to Table 1 for a detailed description of the Kraus operators of the three noise models considered in this work. The noise intensity is tuned by a parameter appearing in the Kraus decomposition, which we indicate with $\gamma$ throughout the paper. Smaller $\gamma$ values correspond to a final quantum state less affected by noise. After the preliminary analysis of the effect of an ideal pure noise, we test our architecture with a more realistic model of quantum noise, which faithfully emulates noise in a superconductive device. To do this, we apply the quantum channels that model the noise of IBMQ backends freely available to the user. The simulation of a backend noise model implemented by Qiskit [39] exploits a combination of thermal relaxation and depolarizing channels. This quantum channel is accessible through a built-in Qiskit feature. In the experiments, we considered the IBM-OSAKA and IBM-SHERBROOKE quantum backend. We refer to Qiskit documentation for a detailed description of realistic models for noise in IBM quantum devices.
II.6 Non-unital channels are essential: a geometrical interpretation
Quantum channels that preserve identity, namely $\mathcal{T}(\mathbb{I})=\mathbb{I}$, are said to be unital. Unital channels are not suitable for reservoir computing, as shown by the following theorem. Here $\mathcal{B}(\mathcal{H})$ is the set of all bounded operators and $d$ is the dimension of the Hilbert space.
Theorem II.4 (Theorem 5 in [40])
Assume there exists $\epsilon>0$ and an operator norm $\norm{\cdot}$ such that the operator $\mathcal{T}(\cdot,u)\colon\mathcal{B}(\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H})$ satisfies $\norm{\mathcal{T}(\cdot,u)}\leq 1-\epsilon$ for any $u\in\mathcal{I}$. Then, the correspondent functional $T$ is constant with $T(u)=\mathbb{I}/d$ if and only if $\mathcal{T}$ is unital for any input, namely $\mathcal{T}(\mathbb{I},u)=\mathbb{I}\quad\forall u\in\mathcal{I}$.
Figure 2: From left to right, the action of a quantum channel on the Bloch sphere. Blue area is the image of each channel. Black and blue dots are the projection of the center of the sphere, respectively, before and after the application of the channel. For clarity, we report simulations for $\gamma=0.6$ and tenfold intensity of fake-OSAKA.
The action of unital channels has a geometrical interpretation in the Bloch sphere representation of qubit state, which allows distinguishing between unitality and non-unitality of a given channel, even without knowing its analytical expression. The set of pure states in the image of a quantum channel $\mathcal{E}$ is called its pure output $\mathrm{PO}\left(\mathcal{E}\right)$, namely
$\mathrm{PO}\left(\mathcal{E}\right)=\left\{\mathcal{E}(\rho),\forall\rho\in\mathcal{S}(\mathcal{H})\right\}\cap\mathcal{P}\,.$
In the Bloch sphere, the pure output is the intersection between the spherical surface of the pure states and the ellipsoid representing the image of the quantum channel. Unital channels fulfill a fundamental symmetry property.
Lemma II.5 (Lemma 4.3 in [41])
The pure output of an unital channel $\mathcal{T}$ is centrally symmetric.
Hence, by looking at the pure output of a quantum channel we are able to determine whether it is unital or not [41]. In particular, the channel is unital if and only if the center of the image ellipsoid coincides with the center of the Bloch sphere. This fact has a trivial interpretation since the center of the Bloch sphere corresponds to the maximally mixed state $\rho=\mathbb{I}/2$. Precisely, we recall the following theorem.
Theorem II.6 (Theorem 4.9 in [41])
Let $\mathcal{E}$ be a qubit channel and $\mathcal{P}$ the set of pure states. One of the following holds:
  1. 1.
    $\mathrm{PO}(\mathcal{E})=\emptyset$, the channel has no pure output;
  2. 2.
    $\mathrm{PO}(\mathcal{E})=\{\xi\}$, the channel has a unique pure output state $\xi$;
  3. 3.
    $\mathrm{PO}(\mathcal{E})=\{\xi,\zeta\}$, the channel has exactly two pure outputs $\xi,\zeta$;
  4. 4.
    $\mathrm{PO}(\mathcal{E})=\mathcal{P}$, all pure states are in the image of $\mathcal{E}$.
In particular, combining Lemma II.5 and Theorem II.6, we can deduce that a channel with exactly two pure outputs is unital if and only if they are antipodal on the Bloch sphere (thus, are orthogonal in the Hilbert space). Figure 2 shows the action of amplitude damping, depolarizing, phase damping and fake-OSAKA channel. As expected, amplitude damping is the sole non-unital ideal noise channel. Moreover, we can observe that the fake-OSAKA channel is non-unital.
III Results
In this work, we fix $N=7$ as the number of qubits. We consider a total timespan of $L=200$. The washout period lasts 10 time steps ($T_{\text{wo}}=10$), while the training and test intervals are 130 and 60 time steps long, respectively ($T_{\text{tr}}=140$). All the experiments are conducted using Qiskit library [39].
III.1 Noise and fading memory
Quantum noise enhances the fading memory of the system by gradually reducing its dependence on inputs from the past. Indeed, the action of the quantum channel modifies the position of each qubit on the Bloch sphere, contributing to the reduction of information over time. To verify whether increasing noise reduces the dependence on previous information stored in the network, we study its effect on the correlation among subsequent measurement outcomes of the same qubit. Precisely, for any measured qubit, we compute $\mathrm{corr}(\overline{z}^{i}_{t},\overline{z}^{i}_{t-k})$, where the vector $\overline{z}^{i}_{t}=\left\{(z_{t})_{s}\right\}_{s=1,\dots,S}$ contains the outcomes of measurements for each shot. Then, we compute the mean correlation over the whole execution by summing over $t$. Figure 3 shows the mean correlation over time for increasing time windows $k$. As expected, higher noise intensity enhances the progressive loss of correlation, aiding the network in avoiding getting stuck in past information. The loss of correlation appears comparable for different noise models. From a theoretical point of view, fading memory is related to the contractivity of the evolution of the echo state network, as discussed in Subsection II.1. We refer to Appendix A for more details on the contractivity of noise models in trace norm. Thus, fading memory alone does not a priori justify the drastic beneficial effect of amplitude damping on the memory capacity of the network and its accuracy in learning tasks.
Figure 3: The decay over time in the mean correlation in subsequent measurements is enhanced by increasing damping intensity. The thickness of lines is proportional to the accuracy in NARMA5 prediction.
Figure 4: Accuracy in reproducing NARMA tasks and memory capacity for increasing intensity of amplitude damping (in blue), phase damping (in grey) and depolarizing channel(in green).
Figure 5: Accuracy in NARMA tasks for noiseless execution, compared with optimal amplitude damping and the realistic noise model of OSAKA backend. Optimal values of noise are $\gamma=0.024,0.03,0.014$ for NARMA2, NARMA5 and NARMA8, respectively. Grey vertical line separates the training and test interval in the output sequence. Red and blue lines represent the target and trained output, respectively.
III.2 Comparison of predicting capabilities under noise models
Ideal noise models.
After confirming that noisy channels ensure fading memory in the system, we investigate their beneficial effect on the predictive capabilities of the architecture using two standard benchmarks for recurrent networks, namely memory capacity and NARMA, presented in detail in Subsection II.4. Figure 4 shows NMSE and MC for increasing intensity of the three noise models considered in our work, amplitude damping, phase damping and depolarizing channel. Remarkably, we observe that amplitude damping drastically enhances memory capacity and accuracy in the prediction of the echo state network, thus confirming the essential role of fading memory in performing time-dependent tasks. Consistently, a higher memory capacity of the network is reflected in better performance in reproducing time-dependent tasks, as can be deduced by comparing results presented in Figure 4. On the other hand, phase damping and depolarizing noise have detrimental effects on both memory capacity and accuracy. As discussed in Subsection II.6, the beneficial effect of amplitude damping is related to its non-unital nature as a quantum channel. Indeed, non-unital channels enrich the dynamics of the reservoir, allowing for a wider exploration of the phase space and, consequently, greater storage of information.
Realistic model
As a further test in preparation for application on a quantum computer, we repeated the experiment by applying a channel that replicates the noise model of some machines provided by IBM. Even if the accuracy in reproducing tasks is slightly deteriorated compared with optimal pure amplitude damping, we still observe a remarkable improvement in learning capabilities, when compared with noiseless executions, as shown by Figure 5. In general, a realistic noise model for real superconductive hardware is the application of several quantum channels, including the three noise models considered in this work. Thus, we may assume that it is strictly contractive since the composition of non-expanding channels and strictly contractive channels is indeed a contractive channel (recall that quantum channels are non-expanding). Moreover, we showed in Subsection II.6 that the realistic noise is non-unital by analyzing its effect on the Bloch sphere.
III.3 Effects of local measuraments
Figure 6: Accuracy in NARMA task with complete measurement of the register. Results for a swarm of 20 executions are plotted in transparency, mean values are in bold in the picture. Running average on a window of ten values of $\gamma$ is performed for smoother visualization.
The emergent properties of a quantum circuit drastically depend on the number of mid-circuit measurements performed on it [35]. This is the case also for the expressivity of the reservoir. Indeed, the beneficial effect of amplitude damping appears stable under increasing noise, at least for small values of $\lambda$ (approximately, $\gamma<0.1$) when performing a complete measurement of the qubit register, as shown in Figure 6. An optimal noise regime, corresponding to which the architecture performs best for any task, emerges for $\gamma\sim 0.02$. This suggests the existence of an optimal critical regime for the network’s expressivity. Interestingly, however, this optimal value of $\gamma$ is followed by a plateau of suboptimal values. This fact sheds light on the feasibility of quantum reservoir computing even in the absence of a reliable quantum error correction procedure. Figure 7 shows how learning capabilities vary under the effect of partial measurements of the qubit register. The plateau of suboptimal values is disrupted in the case of partial measurements of the qubit register. However, the optimal noise regime persists for the same noise intensity, with no deterioration in memory capacity or predictive accuracy.
Figure 7: Accuracy in NARMA task and memory capacity for a different number of measured qubits in the register. Yellow line correspond to complete measurements.
IV Conclusion
We proved the beneficial effect of noise in quantum reservoir computing, as it provides fading memory of the network. In particular, we recognize the noise models characterized by non-unital channels as essential for guaranteeing a drastic improvement in the learning capabilities of a quantum echo state network that employs a register of superconducting qubits as the reservoir. Indeed, we observe that amplitude damping ensures great accuracy in benchmark time-dependent tasks. Improvements in learning are confirmed if the ideal noise model is substituted by a quantum channel emulating the behavior of a realistic superconductive computer, suggesting useful results in real-world applications. Moreover, the beneficial effect of amplitude damping is stable while increasing noise intensity. The latter fact highlights how the quantum reservoir computer can be utilized even without any error correction protocols. As often observed in quantum machine learning, the plateau disappears in the case of local measurements, while the optimal regime persists. As a consequence, by employing effective error control algorithms that allow positioning within this optimal regime, the number of measurements can be drastically reduced, leading to significant savings in execution time.
Acknowledgements
F.M. and E.P. are supported by PRIN-PNRR PhysiComp (nr. grant G53D23006710001). L.N. is partially funded by ENI S.p.A.
Code availability
The code is made available by the authors upon reasonable request.
Competing interests
Authors declares no competing interests.
Appendix A Contractivity of noise channels
In this Appendix, we prove that a noise channel $\mathcal{D}:\mathcal{S}(\mathcal{H})\rightarrow\mathcal{S}(\mathcal{H})$ ensures fading memory by proving that it is strictly convergent in the trace norm. Namely, we show that there exists $r<1$ such that $\norm{D(\rho)-D(\sigma)}_{1}\leq r\norm{\rho-\sigma}_{1}$. For clarity, we write the proof for the single-qubit Hilbert space $\mathcal{H}=\mathbb{C}^{2}$. The extension to the multidimensional case is trivial. Now we propose some common noise channels and evaluate whether they guarantee fading memory or not. In order to do so, it is helpful to Kraus decompose each channel [42], hence write each channel in the form
$\mathcal{D}(\rho)=\sum_{i}K_{i}\rho K^{\dagger}_{i}.$
For the following calculations, it is useful to keep in mind that for a generic Kraus operator $K$ we have
$\begin{split}\norm{K\rho K^{\dagger}}_{1}=K_{ij}\rho_{jl}K^{\dagger}_{li}=k_{i}\delta_{ij}\rho_{jl}k_{l}\delta_{li}=\\ =k_{i}\rho_{ii}k_{i}=k_{i}^{2}\rho_{ii}\end{split}$ (2)
where summation is implied over repeated indices. Remarkably the set of strictly contractive quantum channels is dense in the set of all quantum channels [43]. That is, in the assumption of finite resolution in the observations, we can assume that any channel modeling noise is strictly contractive, thus ensuring fading memory to the system.
Bit flip, bit-phase flip and phase flip.
The Kraus decomposition of the bit flip map is
$\mathcal{D}_{\text{bf}}(\rho)=\sqrt{1-\gamma}\rho+\sqrt{\gamma}X\rho X.$
Notice that the bit-phase flip and the phase flip maps have the same form of the bit flip, but instead of the $X$ Pauli operator, they involve the $Y$ and $Z$ operators, respectively. Hence we have
$\mathcal{D}_{\text{bpf}}(\rho)$ $=\sqrt{1-\gamma}\rho+\sqrt{\gamma}Y\rho Y.$
$\mathcal{D}_{\text{pf}}(\rho)$ $=\sqrt{1-\gamma}\rho+\sqrt{\gamma}Z\rho Z.$
Then, for the sake of conciseness, we only show the case of the bit flip, the other cases being a trivial extension. Since Pauli operators are traceless operators, we can identify
$K_{0}=\sqrt{1-\gamma}\,\mathbb{I}\mbox{\quad and\quad}K_{1}=\sqrt{\gamma}\,X$
By exploiting Equation 2, we have trivially
$\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}\leq 2\left(1-\gamma\right)|\rho_{00}-\sigma_{00}|$
and
$\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}=0.$
Thus, such channels are contractive in trace norm.
Amplitude damping.
The quantum operation that describes energy dissipation is the amplitude damping. Recalling the Kraus decomposition of the amplitude damping channel, we have
$\mathcal{D}_{\text{ad}}(\rho)=K_{0}\rho K_{0}^{\dagger}+K_{1}\rho K_{1}^{\dagger}$
with $K_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}$ and $K_{1}=\begin{pmatrix}0&\sqrt{\gamma}\\ 0&0\end{pmatrix}$. We consider the two terms separately, exploiting Equation 2. Thus, since $\mathrm{tr}(\rho)=\mathrm{tr}(\sigma)=1$, for the first Kraus operator we conclude that
$\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}=|\rho_{00}-\sigma_{00}|\left(1+\sqrt{1-\gamma}\right)\leq$
$\leq r\norm{\rho-\sigma}=2|\rho_{00}-\sigma_{00}|$
with $r=1-\frac{\gamma}{2}<1$. Similarly, for the other term, we get
$\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}\leq r\norm{\rho-\sigma}_{1}$
for $r=\gamma/2$. We conclude that $D$ is strictly contractive, since
$\norm{\mathcal{D}_{\text{ad}}(\rho)-\mathcal{D}_{\text{ad}}(\sigma)}_{1}\leq\norm{K_{0}(\rho-\sigma)K_{0}^{\dagger}}_{1}+$
$+\norm{K_{1}(\rho-\sigma)K_{1}^{\dagger}}_{1}\leq r\norm{\rho-\sigma}_{1}$
with $r=\max\left(1-\frac{\gamma}{2},\,\gamma/2\right)<1$.
Depolarizing channel.
Consider a qubit that has probability $p$ of being depolarized. The action of the depolarizing channel is represented by
$\mathcal{D}_{\text{de}}(\rho)=\frac{p}{2}\mathbb{I}+(1-p)\rho.$
In the operator-sum representation we can write
$\left(1-\frac{3p}{4}\right)\rho+\frac{p}{4}\left(X\rho X+Y\rho Y+Z\rho Z\right)$
or, equivalently, it is more commonly written as
$\left(1-p\right)\rho+\frac{p}{3}\left(X\rho X+Y\rho Y+Z\rho Z\right)$
which means that the state $\rho$ is left alone with probability $1-p$, and each Pauli operator acts with probability $p/3$.
Phase damping.
The loss of information, without loss of energy, is described by the phase damping. We observe that the phase damping channel can be written as
$\mathcal{D}_{pd}(\rho)=\widetilde{K}_{0}\rho\widetilde{K}_{0}+\widetilde{K}_{1}\rho\widetilde{K}_{1}$
with
$\begin{split}\widetilde{K}_{0}&=\matrixquantity(1\hfil&0\\ 0&\sqrt{1-\lambda})\\ \widetilde{K}_{1}&=\matrixquantity(0\hfil&0\\ 0&\sqrt{\lambda})\\ \end{split}$
where the parameter $\lambda$ can be interpreted as the probability of having a photon scattered (without loss of energy) as it travels through a waveguide.
An important thing to notice is that by a unitary recombination of $\widetilde{K}_{0}$ and $\widetilde{K}_{1}$, we can write other two equivalent Kraus operators for the phase damping channel: define $\gamma=\left(1-\sqrt{1-\lambda}\right)/2$, then
$\begin{split}K_{0}&=\sqrt{1-\gamma}\,\mathbb{I}\\ K_{1}&=\sqrt{\gamma}\,X\\ \end{split}.$
Thus, phase damping describes the same operation as the phase flip. Hence it inherits the same contraction and unitality properties.
Thermal relaxation.
Thermal relaxation is equivalent to phase-amplitude damping. The problem is that the Qiskit implementation of the thermal relaxation channel is different from the standard one, hence a short comment is needed. Thermal relaxation, in the Qiskit version, is modeled by the following Kraus decomposition in Qiskit [39]:
$\begin{split}K_{0}&=\sqrt{1-p_{z}-p_{\rm r0}-p_{\rm r1}}\,\mathbb{I}\\ K_{1}&=\sqrt{p_{z}}\,Z\\ K_{2}&=\sqrt{p_{\rm r0}}\,\matrixquantity(1\hfil&0\\ 0&0)\\ K_{3}&=\sqrt{p_{\rm r1}}\,\matrixquantity(0\hfil&0\\ 0&1)\end{split}$
where $p_{z}$ is the probability of phase flip, $p_{\rm r0}$ is the reset probability to state |$\outerproduct{0}{0}$ | and $p_{\rm r1}$ is the reset probability to state |$\outerproduct{1}{1}$ |.